OpenStudy (anonymous):

Simplify the expression. 3 / (5x)^-2 Simplify the expression. 3 / (5x)^-2 @Mathematics

OpenStudy (agreene):

75x^2

OpenStudy (anonymous):

\[\Large \begin{array}{l} \frac{3}{{{{(5x)}^{ - 2}}}}\\ = \frac{3}{{1/{{(5x)}^2}}}\\ = \frac{{3 \cdot {{(5x)}^2}}}{1}\\ = 3 \cdot {5^2} \cdot {x^2}\\ = 75{x^2} \end{array}\]